OpenStudy (anonymous):

Solve the equation for y and then graph it. 16y = −10x can someone please help me with this problem and give me the steps to do it

jimthompson5910 (jim_thompson5910):

16y means 16 times y what operation undoes multiplication?

OpenStudy (anonymous):

Division?

jimthompson5910 (jim_thompson5910):

so you divide both sides by ______ to isolate y

OpenStudy (anonymous):

Okay I do that and then I get stuck with 10x over the 16 I am confused how to isolate the x

jimthompson5910 (jim_thompson5910):

you should have \[\Large y = \frac{-10x}{16}\]

jimthompson5910 (jim_thompson5910):

what is the value of y when x = 0?

OpenStudy (anonymous):

Why is it -10

OpenStudy (anonymous):

And I don't understand the value of y is when the x is 0

jimthompson5910 (jim_thompson5910):

because the start equation is 16y = -10x

jimthompson5910 (jim_thompson5910):

Plug in x = 0 \[\Large y = \frac{-10x}{16}\] \[\Large y = \frac{-10*0}{16}\] \[\Large y = \frac{0}{16}\] \[\Large y = 0\] So when x = 0, y = 0. We now know the point (0,0) lies on the line.

OpenStudy (anonymous):

Okay being dumb here I had it written on the paper wrong

OpenStudy (anonymous):

Can I use a -1 and a positive 1 to make three points by the same operation

jimthompson5910 (jim_thompson5910):

you can use any x value you want (as long as its in the domain)

OpenStudy (anonymous):

How do you give a medal so I can give one to you

jimthompson5910 (jim_thompson5910):

you click "best response" button

jimthompson5910 (jim_thompson5910):

what 3 points do you have?

OpenStudy (anonymous):

I used -1,0,1 and I would like to know if the 16 falls off when you multiply the top

jimthompson5910 (jim_thompson5910):

multiply the top by what?

OpenStudy (anonymous):

-1,0,1

jimthompson5910 (jim_thompson5910):

well in the case of x = 0, you'll have 0/16 = 0

OpenStudy (anonymous):

I got that now its the -1, 1 that confuses me over the 16

jimthompson5910 (jim_thompson5910):

with x = 1, you'll have \[\Large y = \frac{-10x}{16}\] \[\Large y = \frac{-10*1}{16}\] \[\Large y = \frac{-10}{16}\] \[\Large y = -\frac{5}{8}\] \[\Large y = -0.625\]

jimthompson5910 (jim_thompson5910):

if x = -1, then... \[\Large y = \frac{-10x}{16}\] \[\Large y = \frac{-10*(-1)}{16}\] \[\Large y = \frac{10}{16}\] \[\Large y = \frac{5}{8}\] \[\Large y = 0.625\]

OpenStudy (anonymous):

Okay so I used -8,0,8 and I got x=-8 y=5 first point x=0 y=0 for the second point and x=8 y=-5 for the third point I did this by dividing 16 into 80

jimthompson5910 (jim_thompson5910):

those values work

OpenStudy (anonymous):

Yay Jim Thompson you are my hero

jimthompson5910 (jim_thompson5910):

I'm glad it's all clicking now

OpenStudy (anonymous):

I might be asking you more questions lol you are a good teacher

OpenStudy (anonymous):

The linear equation L = −0.04t + 10 estimates the length L (in inches) of a pencil after it has been inserted into a sharpener and the handle turned a total of t times. Use the graph to estimate the length of the pencil after 250 turns of the handle

OpenStudy (anonymous):

The linear equation L = −0.02t + 9 estimates the length L (in inches) of a pencil after it has been inserted into a sharpener and the handle turned a total of t times. Use the graph to estimate the length of the pencil after 250 turns of the handle

OpenStudy (anonymous):

so I did -0.02(250)+9=4 is that right?

jimthompson5910 (jim_thompson5910):

You should do L = -0.04t + 10 L = -0.04*(250) + 10 ... replace t with 250 L = 0 So after 250 turns, the pencil is basically gone (turned into dust probably)

jimthompson5910 (jim_thompson5910):

On the graph, you'll have the point (250,0) which is the x-intercept. The point where the graph crosses the x axis.